Presentation is loading. Please wait.

Presentation is loading. Please wait.

MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 12, 2010.

Similar presentations

Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 12, 2010."— Presentation transcript:

1 MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 12, 2010

2 April 27, 2009Commissioning Needs-Soler2 Commissioning Needs Decay Solenoid —As you all know the Linde refrigerator developed a blockage fault —Linde engineer worked 23-26 March: changed charcoal in gas panel adsorber, heated it out, purged it and refilled compressor and no leak was found. —But when the compressor was isolated from cold box and ran system with cold trap in circuit the N 2 level rose from 50 ppm by 2-3 ppm/hour. —Working hypothesis: changing charcoal had no effect so there might be a leaking compressor shaft seal and might need replacing —The collaboration meeting and Easter got in the way of a speedier resolution —Linde will send their engineer, Matthias Hoberich, and he arrives Tuesday night (so should be ready for work Wed-Sat 14-17 April). —HPC should receive shaft seal today and will also send technician on Wednesday for installation. They will be working with Mike Courthold and Matt Hills this week to try and solve problem. —Once refrigerator is fixed then it will take ~10 days to cool down to operating temperature. —All going well, we will not have the Decay Solenoid ready for data taking until the very end of April or beginning of May

3 April 27, 2009Commissioning Needs-Soler3 Commissioning Needs Decay Solenoid (cont.) —Another problem with PTR90 Pirani/Penning vacuum gauges (giving false readings 4E-3 mbar) – 9 April o Examined with Leybold rep but they were not significantly contaminated o These were cleaned and now they work, but do not really understand why o Purchase four more PTR90 heads for spares (£703 each) —Scroll pumps have never been serviced o DS south side pump struggling to pull vacuum below 3mbar – removed for servicing o DS north side pump can cope below 2E-3 mbar, so that pumping stack now in full use and pumping the DS! o Transfer line pump removed and replaced DS south side pump but also needs servicing and can only go down to 0.2 mbar

4 April 27, 2009Commissioning Needs-Soler4 Commissioning Needs Target —Investigation of R78 target: see collaboration meeting slides —Target in ISIS has a total of 161K actuations and working well Detectors —Luminosity Monitor o LM stand was bolted to ISIS floor during April shutdown o Dedicated machine physics run on Friday 16 April to determine dead time of LM (change gate width as a function of beamloss and measure rate) —Tracker o Problems with cryostat: due to cold weather over Christmas, seal was damaged and developed leak. o Alan Bross at RAL this weeek to fix problem. —CKOV,TOF0, TOF1, TOF2, KL: I would need experts to start these up in preparation of run. o TOF0,1,2 would need new calibration? Beam Magnets —Willie installed new pressure valves in D2, Q4-Q9 warm magnets. —Flow rate is now 60-67 l/min in quadrupoles Q4-Q9 and dipole D2 (used to be around 30-32 l/min). —Q1-Q3 and D1 remain with flow rate ~30 l/min —Magnet polarity labelling scheme from Henry has been adopted

5 April 27, 2009Commissioning Needs-Soler5 Preparation for run Machine development period: 12-19 April —We have only requested one shift (18:00-24:00) during this period, on Friday 16 April for Luminosity Monitor rate vs gate width studies (change LM gate from 10-40 ns and beamloss 0.5- 2.0V) Programme ISIS run: 20 April – 28 May —While decay solenoid not operational (20 April – first or second week May?) o Perform new TOF0,1,2 calibration with blown up defocused beam to try and cover all counters – would need beam parameters from Marco o Proton absorber tests – move different thickness of absorber to compare particle rate vs absorber thickness —While decay solenoid operational (First or second week May-26 May): o Perform target scan: delay target vs depth (ie acceleration) to try and flatten spill and reduce pileup and beam overlaps. o Work our way through Matrix of emittance (4 , 6 , 10  ) vs momentum (160, 200, 240 MeV/c) with all detectors on. o Statistics for each point? ~1Million muons?

6 April 27, 2009Commissioning Needs-Soler6 Preparation for run ISIS shutdown: 29 May-13 June Next machine development period 14-21 June —Request further activation run: 16 hours with 4V beamloss? ISIS users run 22 June – 12 August —Continue with Step I programme of emittance measurements vs momentum matrix. —Last ISIS run before long shutdown 13 August 2010-February 2011

Download ppt "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 12, 2010."

Similar presentations

Ads by Google